# Lookup multivariate Kolmogorov-Arnold Networks

### v1

@@ F.5 ImageNet
The standard data preparation pipeline introduced by AlexNetKrizhevsky et al. [2012] involves first resizing an image to $256$ pixels along the smallest dimension, then performing a random crop of $224\times 224$ pixels during training, and a center crop of $224\times 224$ pixels during validation.
We mimic this procedure by first resizing the image to $81*256/224\approx 93$ pixels across smallest dimension, and then performing random or center crops of $81\times 81$ pixels.
Next, we use the following augmentational pipeline:
@@ F.6 Comparison with FastKAN
The data was split properly into train, validation, and test subsets, while the original script employed only train-validation split. We use the cosine (without restarts) learning rate scheduler Loshchilov and Hutter [2016] instead of the exponential decay one of the original script. Finally, the normalization was performed with real values of mean and std for the CIFAR-10 dataset, instead of the dummy $0.5$ values of the original script.

### v2

@@ F.5 ImageNet
The standard data preparation pipeline introduced by AlexNet[10] involves first resizing an image to $256$ pixels along the smallest dimension, then performing a random crop of $224\times 224$ pixels during training, and a center crop of $224\times 224$ pixels during validation.
We mimic this procedure by first resizing the image to $81*256/224\approx 93$ pixels across the smallest dimension, and then performing random or center crops of $81\times 81$ pixels.
Next, we use the following augmentation pipeline:
@@ F.6 Comparison with FastKAN
The data was split properly into train, validation, and test subsets, while the original script employed only a train-validation split. We use the cosine (without restarts) learning rate scheduler [66] instead of the exponential decay of the original script. Finally, the normalization was performed with true values of the mean and standard deviation for the CIFAR-10 dataset, instead of the placeholder $0.5$ values of the original script.